OpenStudy (lalaland_lauren):

Given the functions f(x) = 6x + 11 and g(x) = x + 6, which of the following functions represents f[g(x)] correctly? f[g(x)] = 6x + 47 f[g(x)] = 6x + 17 f[g(x)] = 6x2 + 47 f[g(x)] = 6x2 + 17

OpenStudy (lalaland_lauren):

Can someone help explain this to me?

imqwerty (imqwerty):

f(g(x)) means that you have to put g(x) as the value of x in f(x)

OpenStudy (campbell_st):

just replace x with x + 6 so f(x +6) = 5(x + 6) + 11 then just simplify

OpenStudy (campbell_st):

oops should be f(x + 6) = 6(x + 6) + 11 sorry for the typo

imqwerty (imqwerty):

so u'll have - f(g(x)) = 6(g(x))+11 f(g(x))=6(x+6)+11 f(g(x))=6x+36+11 f(g(x))=6x+47

OpenStudy (lalaland_lauren):

so it'd be a

imqwerty (imqwerty):

yes

OpenStudy (lalaland_lauren):

okay, could you help me with some more?

imqwerty (imqwerty):

yes

OpenStudy (lalaland_lauren):

Given the following functions f(x) and g(x), solve f[g(6)] and select the correct answer below: f(x) = 6x + 12 g(x) = x − 8 −96 0 24 48

OpenStudy (lalaland_lauren):

hello?

OpenStudy (campbell_st):

just find the value of g(6) = 6 - 8 then substitute the answer into f(x)

OpenStudy (lalaland_lauren):

okay so g(6) = -2 then I divide?

OpenStudy (campbell_st):

then you substitute f(-2) = 6(-2) + 12

OpenStudy (lalaland_lauren):

oh okayyy

OpenStudy (lalaland_lauren):

you're a huge help lol

OpenStudy (lalaland_lauren):

can you help me with two more? @campbell_st